Grade 12th passMechanics

X-rays of wavelength 0.122nm are formed to reflect in the second order from the face of a lithium flouride Crystal at a Bragg angle of 28.1*. Calculate the distance between adjacent crystal planes.

Salman khan
8 Years agoGrade 12th pass

1 Answer

Phanindra
8 Years ago
According to Braggs equation...n × (wave length) = 2d sin(theta)..Here n = 2..Wave length = 0.122 × 10 ^-9 m..d = ?Theta = 28.1 so... d = (2 × 0.122 × 10 ^-9) ÷ (2 × sine28.1)... d = 0.7 × 10 ^-9 m..d = 7 × 10^ -10 m ...So...d = 7 Angstroms..